OpenStudy (anonymous):

Solve the equation. square root(6x +8) = square root(7x-6)

OpenStudy (anonymous):

\[\sqrt{6x+8}=\sqrt{7x-6}\]

OpenStudy (anonymous):

solve \(6x+8=7x-6\)

OpenStudy (anonymous):

14=x...think I am getting tired, lol

OpenStudy (anonymous):

\[\sqrt{6x+8}=\sqrt{7x-6}=>(\sqrt{6x+8})^{2}=(\sqrt{7x-6})^{2}\](squaring both sides) 6x+8=7x-6 solve it @maliperez

OpenStudy (anonymous):

to get rid of the square root sign, I have to square the problem, correct? so the answer would be 14...am I close?

OpenStudy (anonymous):

yup u r rite nw chech if LHS=RHS by putting the value of x in both sisdes @maliperez

OpenStudy (anonymous):

it equals 92 on both sides, thanks...I think it is time for bed, I can hardly concentrate, lol

OpenStudy (anonymous):

ok gud n8 frnd bt itz morning in our country :-) @maliperez

OpenStudy (anonymous):

its midnight here, and I have to be back up at 5am..so have a wonderful day with lots of sunshine :)

OpenStudy (anonymous):

thnq frnd.......have a wonderful and dreamful night :-)@maliperez

OpenStudy (anonymous):

yeah \(14=x\) is right